OpenStudy (just_one_last_goodbye):

Can someone help me with Fractions? :/ I know it's stupid but I was never good at fractions

OpenStudy (just_one_last_goodbye):

1/3 = ?/12

OpenStudy (just_one_last_goodbye):

^ how do I make that true?

OpenStudy (just_one_last_goodbye):

I have some guesses of what the answer might be :/

OpenStudy (anonymous):

well :/ how do you get 3 to 12?

OpenStudy (just_one_last_goodbye):

3 times 4 :)

OpenStudy (just_one_last_goodbye):

so is the answer 4?

Parth (parthkohli):

try making two cakes of the same sizes. in one cake, cut it into three pieces and eat one. in the other cake, if you cut it into twelve pieces, how many pieces will you eat to eat the same amount of the cake?

OpenStudy (just_one_last_goodbye):

well 3 - 1 = 2 so... 12 - 10 =2

OpenStudy (just_one_last_goodbye):

is that right? ._.

OpenStudy (just_one_last_goodbye):

im sorry ;-; Im just in pain and need to go to the hospital and cant go till I finish all my school ;-;

OpenStudy (anonymous):

okay do it this way 1/3 > 2/6 >?

OpenStudy (just_one_last_goodbye):

what do you mean Reign?

OpenStudy (triciaal):

please stay positive. It is not stupid if you are not good at fractions. not everything is easy for everyone. you play the piano and the drums very well not everyone plays an instrument

OpenStudy (just_one_last_goodbye):

thank you @triciaal

OpenStudy (anonymous):

okay 1/3 equals 2/6 right?

OpenStudy (just_one_last_goodbye):

The answer is 4 ._.

OpenStudy (triciaal):

to make equivalent fractions multiply the lowest fraction by whole numbers expressed as 1

OpenStudy (anonymous):

yups :P

OpenStudy (just_one_last_goodbye):

I asked if it was 4 at the beginning and no one said yes or no :/ so thought it was wrong

OpenStudy (triciaal):